OpenStudy (anonymous):

i REALLY need help on this, please. Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system. x + 6y = 1 2x + 5y = 9 i REALLY need help on this, please. Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system. x + 6y = 1 2x + 5y = 9 @Mathematics

OpenStudy (beefarm):

Using one of the equations, express one of the variables in terms of the other. Then plug that expression into the other equation, solve for the first variable then plug that value into either of the equations to solve for the other variable. In this particular case: x+6y =1 x=1-6y //this is the expression for one variable in terms of the other variable 2(1-6y)+5y = 9//we now plug that expression into the other equation 2-12y+5y = 9 2-7y=9 -7y=7 y=-1//we have solved for one variable x = 1-6*(-1)//now we plug the value of the first variable into one of the equations with both variables and solve for the next variable x = 7

OpenStudy (anonymous):

ohmygod, thank you sooooo much.